a right triangle has one 14 degree angle, what are the other angles?
Answer by cleomenius(959)

180 - 90 = 90 degrees,'
90 - 14 degrees = 76 degrees.
76 + 14 + 90 = 180 degrees.
This does check.
Cleomenius.